OpenStudy (anonymous):

Find the percent of the number. 2.3% of 80 A. 1.84 B. 18.4 C. 184 D. 1840

rebeccaxhawaii (rebeccaxhawaii):

well lets use common sense .. we can cancel out c and d because its more than 80

rebeccaxhawaii (rebeccaxhawaii):

and 18.4 is way more that 2.3% of 80

rebeccaxhawaii (rebeccaxhawaii):

so that leaves you with only one answer

rebeccaxhawaii (rebeccaxhawaii):

but heres an example for references for further questions

rebeccaxhawaii (rebeccaxhawaii):

To determine the percent of a number do the following steps: Multiply the number by the percent (e.g. 87 * 68 = 5916) Divide the answer by 100 (Move decimal point two places to the left) (e.g. 5916/100 = 59.16) Round to the desired precision (e.g. 59.16 rounded to the nearest whole number = 59)

OpenStudy (anonymous):

Just do \[\large \sf .03 \times 80\]

OpenStudy (anonymous):

Like lol

OpenStudy (anonymous):

Can You Help With Another? Find the percent of the number. 113% of 140 A. 15,820 B. 1.582 C. 1582 D. 158.2

rebeccaxhawaii (rebeccaxhawaii):

so take 13% out of 140 and add that to 140

rebeccaxhawaii (rebeccaxhawaii):

so that rules out a,b and c

OpenStudy (anonymous):

32 is 80% of what number? A. 0.4 B. 4 C. 25.6 D. 40

rebeccaxhawaii (rebeccaxhawaii):

ok you need to learn to do it on your own just use common sense

rebeccaxhawaii (rebeccaxhawaii):

there is only one logical answer for that one

rebeccaxhawaii (rebeccaxhawaii):

it needs to be greater that 32 because that is 80% of a number which can only be one answer

OpenStudy (anonymous):

Would It Be C?

rebeccaxhawaii (rebeccaxhawaii):

pls try to figure it out on your own now ive helped you on multiple questions and you seem to be giving 0 effort other than copying and pasting your questions

rebeccaxhawaii (rebeccaxhawaii):

how would it be c

rebeccaxhawaii (rebeccaxhawaii):

if 32 is 80% then you would have a greater number than 32 because 32 is ONLY 80% of the number

OpenStudy (anonymous):

Im Sorry >.< I Don't Know If Im Doing It Right This Is How I Did It 80% = .8 32 × .8 = 25.6 Am I Doing It Wrong?